Project

General

Profile

Ops/Comm Scheduling meeting

Location : CR25 in R1 @ 15:30 pm on Tuesday 7th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

1. Agenda

Actions from last meeting
  • Generate appropriate table to store absorber pressure and material information (Janusz) : (almost done)
  • DS readback current requires calibration (DL) : stands until close to powerup
  • Henry, Craig, Paul and Pierrick to look into the water detectors in the trench : Still being worked on. Stands
  • Document 'magnet off' buttons (Josef), when complete send to Andy & DC's. Must be a standalone document.
  • Compressor failure alarm for ISIS monitoring (Pierrick)
  • Update electrical work document (Steve G/Andy G)
  • Press the US team for a work plan (DC)
  • Write a document on what to do if the compressors or the water fail (Josef, John, DC, Pierrick)
  • Get a record of the measurements of PRY movements that have been made to the engineers to let them think about it
  • Discuss dedicated movement measurements. (Victoria/Pierrick)
  • Check that PRY movement monitor measurements are in the archiver (Victoria).
  • Verify the sensors are operating as expected (Colin/John/Pierrick)

Ongoing work

  • SS QD/QP commissioning
    • DCCT readout on rampdown
    • Lakeshore studies (Pierrick, requires SC permit & possible shorting at link box) -- Wednesday afternoon & all day Friday
  • Electrical work
  • Water system work
    • DSA search & secure this week
    • Alarm limits on compressor output water temperature
  • FC cooldown
  • Target fibre work
  • ODH alarm issues
  • PRY movement issues
  • Planning for July running
    • Expert startup date :
    • Issues from last run
      • DATE crash on first run of the day
      • Proton absorber sticking
      • GVA1 (?)
      • DS Fridge instability
      • Nagios/Heart beat alarms
      • Q4 bounce
      • Q6 leak
      • PPS Door lock (bouncing)
      • CKOV monitoring
      • HV channels (?)
      • Sick shifters (?)
      • Trench water leak / Q5 water leak
      • Compactor file corruption
      • miceraid5 disk full
      • Backups not working for a fortnight
    • What needs to be ready to run?
      • SSH lockout procedure
      • Cooling water system
      • Controls and Monitoring including SSU/SSD/FC : Expert panel readiness
      • Beam/Target tuning session
      • Add items here...
    • BLOC training, 7th & 8th June
  • AOB
  • DONM : 14/6/2016

Minutes

Present : P. Hanlet, M. Courthold, V. Bayliss, C. Rogers, M. Tucker, J. Bohm, C. Macwaters, V. Blackmore, C. Whyte, A. Gallagher, J. Govans, K. Long, E. Overton, A. Bross, E. Overton, S. Boyd, J. Cobb

Actions from last meeting
  • Generate appropriate table to store absorber pressure and material information (Janusz) : Will be addressed on Friday
  • DS readback current requires calibration (DL) : Asked Daresbury when they can schedule it. Possibly next week depending on workload.
  • Henry, Craig, Paul and Pierrick to look into the water detectors in the trench : Still being worked on. Stands
    • Decided to remove this option and put a webcam in the trench. Water sensors for the DSA will be developed for August installation.
  • Document 'magnet off' buttons (Josef), when complete send to Andy & DC's. Must be a standalone document.
    • Buttons all installed and need to be tested. Once tested document will be written and distributed.
  • Compressor failure alarm for ISIS monitoring (Pierrick) Done. Should add one for the focus coil.
  • Update electrical work document (Steve G/Andy G) : Stands, but DL team is aware.
  • Press the US team for a work plan (DC) : Done. Require a ramp-up plan from Jaroslaw.
    • Action : get a ramp-up plan from Jaroslaw. (Steve, Chris)
    • Josef wants to start ramping the FC next week to see how it behaves.
    • Action : Josef send FC plan to Colin/DC.
  • Write a document on what to do if the compressors or the water fail (Josef, John, DC, Pierrick)
    • A document has been written with emergency contact details. A schematic drawing is being prepared with an action plan.
  • Hall probes in tracker are being unstable. Melissa has been asked to look into it.
    • Action : Ops Coord/DC to chase Melissa.
  • Get a record of the measurements of PRY movements that have been made to Jason to let them think about it. Done. Engineers
    don't know if the movement seen is acceptable. All we can do is make sure that the installation follows Steve Plates instructions.
    Alan has asked Steve to give us some numbers for 100,200 and 240 A as a guide for the expected deflection of the PRY so we know
    what to expect. Colin: can we put some gauges which measure lift to ensure that the PRY is not rotating. Our current gauges measure
    translation - we also should have some for rotation around one of the legs. Alan: should have a wire sensor on both upper and lower
    East and West endplates.
    • Action : look at putting wire gauges onto end plates (John/Pierrick)
  • Discuss dedicated movement measurements. (Victoria/Pierrick) : Done
  • Check that PRY movement monitor measurements are in the archiver (Pierrick) : Stands
  • Verify the sensors are operating as expected (Colin/John/Pierrick) : Done.

Ongoing work

  • SS QD/QP commissioning
    • Fix functionality of new resistor packs
      • This will happen next week when Trevor comes down.
    • Lakeshore studies (Pierrick, requires SC permit & possible shorting at link box) -- Wednesday afternoon and Monday have been reserved for this.
  • Electrical work
    • Andy : needs to look at the problem with the dump resistors. Also will want to do the safety checks on the FC before current
      can be put on FC. Water circuits in the trench have a new control panel.
  • Water system work
    • everything is there. Electrics are all in. We are working on one compressor at the time. Chiller is running but John is worried
      about pump pressure.
  • DSA search & secure this week
    • Test of conventional magnets to be done on Thursday, followed by return of the DSA to ISIS.
  • Alarm limits on compressor output water temperature
    • Josef talked to Pierrick about resetting the limits to stop the alarms going off. We are far from any danger level. Josef suggests we set the limits of the output water temperature to 45 degrees or so.
  • FC cooldown
    • Coming down. It's in the 80K range now. Bolts on cryocoolers are a bit loose. Josef hand-tightened them to see if there was any
      change on the shield. Expects we will be cold by Monday but we need to wait for Daresbury team to verify all the safety aspects. We
      probably want a dewer here by Tuesday.
  • Target fibre work : All done.
  • ODH alarm issues : All done.
  • PRY movement issues : Already discussed.
  • Planning for July running
    • Expert startup date : : Monday 27th is expert startup
    • Issues from last run
      • DATE crash on first run of the day : Done
      • Proton absorber sticking : Hasn't been reproduced lately - shifters will need to monitor it.
      • GVA1 (?) : Do we pull it out? Yes - John will do it.
      • DS Fridge instability : Has been investigated by Mark. Monitor. Will be tested on Wednesday.
      • Nagios/Heart beat alarms : Done.
      • Q4 bounce : Possibly a wrong regulator board. Daresbury will look at this.
      • Q6 leak : wouldn't touch it - too fragile.
      • PPS Door lock (bouncing) : train people not to bang doors.
      • CKOV monitoring : Should now be operational. Will be checked at expert run.
      • HV channels (?) : Some HV channels are reporting as -1 in the 527. Check at expert run.
      • Sick shifters (?) : Don't come in if dying. Tell MOM if sick.
      • Trench water leak / Q5 water leak : Q5 has been fixed.
      • Compactor file corruption : Looks like two processes were running simultaneously which shouldn't happen. We just have to monitor this.
      • miceraid5 disk full : Paolo has updated Nagios so it will keep an eye on this. We are down to 250 GB so we should do a preemptive
        cleanup. Get in touch with Yordan on self-cleanup.
      • Backups not working for a fortnight : Refer to Paolo to see if we need an alarm for backups.
        • From Paolo : there was a stale mountpoint, because probably something has been moved
          around in the rack room. All the experts receive a mail once per week
          with the status of the micenet backups, I will consider a nagios alarm
          on that,
    • What needs to be ready to run?
      • SSH lockout procedure : tested and working. Need an implementation - (Discuss with Pierrick)
      • Cooling water system
      • Controls and Monitoring including SSU/SSD/FC : Done
      • Expert panel readiness : Done
      • Beam/Target tuning session : Ed tested target on new ISIS timing so new timing settings have been defined (40/64 Hz).
      • Paper copies of documentation should be checked so that it is up-to-date (Me)
      • Shifter training procedures for cooling channels : form action committee
      • Add items here...
    • BLOC training, 7th & 8th June
  • AOB
  • DONM : 14/6/2016

2016-06-06-Workplan.xlsx (11.1 KB) Blackmore, Victoria, 06 June 2016 11:10